# Mapping the Potential of Explainable Artificial Intelligence (XAI) for Fairness Along the AI Lifecycle

### v2

@@ Abstract.
The widespread use of artificial intelligence (AI) systems across various domains is increasingly highlighting issues related to algorithmic fairness, especially in high-stakes scenarios. Thus, critical considerations of how fairness in AI systems might be improved, and what measures are available to aid this process, are overdue. Many researchers and policymakers see explainable AI (XAI) as a promising way to increase fairness in AI systems. However, there is a wide variety of XAI methods and fairness conceptions expressing different desiderata, and the precise connections between XAI and fairness remain largely nebulous. Besides, different measures to increase algorithmic fairness might be applicable at different points throughout an AI system’s lifecycle. Yet, there currently is no coherent mapping of fairness desiderata along the AI lifecycle. In this paper, we set out to bridge both these gaps: We distill eight fairness desiderata, map them along the AI lifecycle, and discuss how XAI could help address each of them. We hope to provide orientation for practical applications and to inspire XAI research specifically focused on these fairness desiderata.
1. Introduction
@@
The emergence and widespread use of artificial intelligence (AI) systems across various sectors and domains is increasingly shifting attention from considerations of mere performance to considerations about algorithmic fairness. This is particularly relevant for systems employed in high-stake scenarios and especially pressing in contexts prone to harmful societal biases (Barocas and Selbst, 2016; Ntoutsi et al., 2020). Thus, critical considerations of how fairness in AI systems might be improved, and what measures are available to aid this process, are overdue. In the literature, there is a common recognition that fairness in AI systems demands various perspectives and measures (e.g., (De-Arteaga et al., 2022; Starke et al., 2022a; Aler Tubella et al., 2023)). In this paper, we set out to integrate two strategies that have been suggested: First, a growing community of researchers proposes explainable AI (XAI) as a versatile and powerful tool to combat unfairness (Deck et al., 2023). Second, others have focused on the AI lifecycle and tried to determine where fairness issues originate (e.g., (Fazelpour and Danks, 2021; Agarwal and Agarwal, 2023)). Their hope is that, once identified, issues with fairness can be mitigated by taking appropriate steps in the relevant phase within the lifecycle.
@@
While both these strategies seem intuitively promising, neither currently presents a satisfactory and comprehensive picture. A primary reason why neither approach currently fulfills their potential is, we take it, that there are various types and kinds of fairness discussed in the literature, yielding different fairness desiderata. Differentiating between these desiderata is crucial to gain an overall picture of which measures are most promising to address fairness in which contexts. Without such differentiation, the utility of XAI is not as straightforward as commonly claimed. Accordingly, there is a need for more clarity about how, exactly, XAI contributes to which fairness desideratum (Deck et al., 2023). Existing discussions on this matter suffer, by and large, from both a too narrow conception of XAI and a mostly under-specified notion of fairness. Similarly, existing attempts to map measures for achieving fairness onto the AI lifecycle remain crucially incomplete; for they have limited their attention to only a subset of the relevant fairness desiderata. To overcome these limitations, we distill eight fairness desiderata, map them along the AI lifecycle, and discuss how XAI can help address each of them.
@@
For the purposes of this paper, we will focus specifically on how XAI can be utilized to improve AI systems’ fairness throughout their lifecycle. With respect to this question, we aim to develop a holistic account based on the fairness desiderata we distill. While our proposal is not intended as a guideline, we believe it will stimulate scientific discussions about and practical applications of fairness measures; and as such will be a valuable starting point to explore fairness opportunities for researchers, developers, and regulators alike.
@@
We begin by introducing preliminaries and diagnosing the problem in Section 2. Importantly, we do not limit our understanding of algorithmic fairness to computational perspectives. In Section 3 we propose eight fairness desiderata: fairness understanding, data fairness, formal fairness, perceived fairness, fairness with human oversight, empowering fairness, long-term fairness, and informational fairness (contribution #1). We offer a mapping of our fairness desiderata along the different stages within the AI lifecycle and suggest that each fairness desideratum affords a different entry point for taking measures to improve fairness. This closes the first gap in the current literature as it highlights where in the lifecycle different fairness desiderata become especially relevant (contribution #2). Utilizing this mapping, we discuss how XAI can be leveraged to address the different fairness desiderata at the respective points throughout the AI lifecycle. This closes the second gap in the current literature as it allows to systematically examine the potential of XAI to address algorithmic fairness in different circumstances (contribution #3). To illustrate the utility of our approach, we discuss its application to the COMPAS case (see (Angwin et al., 2022)) in Section 4. Before closing, we point to some avenues for future research in Section 5.
2. Background
@@
In this article, we contribute to closing two research gaps: 1) considerations of fairness along the AI lifecycle are incomplete, and 2) it is unclear how exactly XAI can help in fostering fairness. Before closing these gaps, we retrace contributions and debates of prior works.
2.1. Algorithmic Fairness
@@
The debate on algorithmic fairness draws inspiration from various disciplines. Several scandals surrounding AI systems that disadvantage marginalized groups (Angwin et al., 2022; Dastin, 2022; Fuster et al., 2022) have shattered early hopes put into the “neutrality” of AI (Barocas and Selbst, 2016). In reaction, scholars from computer science, philosophy, social science, law, and psychology have been engaging in debates on what it means for algorithmic decision-making to be fair (see (Mulligan et al., 2019) for interdisciplinary perspectives).
@@
The technical debate on algorithmic fairness primarily focuses on formal measures of fairness (Barocas et al., 2019; Verma and Rubin, 2018; Castelnovo et al., 2022). After it has proven unhelpful to remove information about membership in marginalized groups from training data (“fairness through unawareness” (Dwork et al., 2012)), most approaches from the field of computer science rely on comparing the outcomes for people from different groups. An important finding is that the different formal measures for fairness can, in most cases, not be fulfilled simultaneously (Kleinberg et al., 2017; Defrance and de Bie, 2023). Thus, (formally) fair AI systems require crucial design choices about which formal measures to apply in which context.
@@
To aid such decisions, philosophers have connected formal measures to different philosophical theories (Binns, 2018; Lee et al., 2021; Khan et al., 2022). In general, “fairness” is a normative concept—and debates about fairness can be understood as a way to discuss what is morally right or wrong in a given case (see (Mulligan et al., 2019)). More specifically, fairness can be understood as an issue of justice, non-discrimination, or equality (Binns, 2018), which connects it to numerous strands of philosophical discussion.
@@
Beyond that, social scientists have pointed out shortcomings of existing formal methods, e.g., problems in detecting issues of intersectionality (Crenshaw, 1991) and their reliance on constructed categories like race and gender (Hanna et al., 2020). Legal scholars are primarily interested in how legislation on discrimination can be applied to algorithmic decision-making and whether additional regulation is required (Wachter et al., 2021; Barocas and Selbst, 2016). Psychological research is generally interested in whether algorithmic decision-making is perceived as fair (Starke et al., 2022a; Lee, 2018). For example, Colquitt and Rodell (2015) distinguish four different psychological dimensions of fairness: Whether the outcome is fair (viz., distributive justice), whether the process that leads to the outcome is fair (viz., procedural justice), whether the information about the decision is communicated truthfully and thoroughly (viz., informational justice), and whether individuals are treated respectfully (viz., interpersonal fairness).
@@
Overall, algorithmic fairness can be addressed from several different angles—some of which are complimentary, some of which are contradictory. This stresses the need to be clear about the specific desideratum behind any attempt to improve fairness.
2.2. Considerations of Fairness Along the AI Lifecycle Remain Incomplete
@@
A common strategy to determine potential entry points to address fairness issues involves examining each step of the AI lifecycle. There have been several proposals of prototypical AI lifecycles (e.g., (Wirth and Hipp, 2000; Fayyad et al., 1996; Wang et al., 2017; Quemy, 2020; Kühl et al., 2021)), each with slightly different stages or adapted to different application areas. The AI lifecycle used in this article is a combination of the proposals by Quemy (2020) and Wang et al. (2017) involving the following stages: 1) problem formulation, 2) data collection, 3) data analysis, 4) feature selection, 5) model construction, 6) model evaluation, 7) deployment, and 8) inference and usage (see Figure 1).
Figure 1. The AI Lifecycle we use, combined from Quemy (Quemy, 2020) and Wang et al. (Wang et al., 2017).
@@
First, problem formulation involves abstraction and formulation of the problem, which is to be solved using AI. Afterward, data collection aims to establish a representative data set suitable to develop an AI system able to solve the defined problem. Data analysis includes descriptive statistics to understand the characteristics of the data at hand and data pre-processing to prepare the data for further operations. Subsequently, in the feature selection stage, features are excluded, transformed, or aggregated for effective and efficient handling in the training process. Model construction, then, includes selection of the training algorithm and the training itself. Iteratively, performance objectives are tested during model evaluation and optimized through adjustments to the algorithm, e.g., in the form of parameter tuning. Deployment refers to the integration of the AI system into a productive environment once the model achieves sufficient performance. Finally, inference & usage describes concrete output generated by the AI system and its impact on the surroundings, such as the business context, society, and environment. Note that these stages are often not entered sequentially but leave room for iterations and loops between stages.
@@
Much research has focused on what types of bias can emerge in or affect different steps of the AI lifecycle (e.g., (De-Arteaga et al., 2022; Agarwal and Agarwal, 2023; Lee and Singh, 2021; Singh et al., 2022)). For example, Singh et al. (2022) take the CRISP-DM model (Wirth and Hipp, 2000) and review the types of bias that can occur at different stages of this model. However, these considerations remain incomplete, as most of these works have a strong technical focus and deal only with formal notions of fairness, e.g., by providing guidance for choosing fairness metrics or debating the role of potential fairness-utility tradeoffs. Against this background, it is important that we develop a view that also maps other conceptions of fairness into the lifecycle.
2.3. Considerations of XAI for Fairness Remain Incomplete
@@
To amend the lack of understanding of AI-based systems, their reasoning processes, and their outputs, the research field of XAI emerged in recent years (Barredo Arrieta et al., 2020; Páez, 2019; Chazette et al., 2021; Langer et al., 2021b; Köhl et al., 2019). Generally, the goal of XAI is to provide a stakeholder with information about some aspect of an AI system to facilitate their understanding of this aspect (Chazette et al., 2021). Understanding (some aspect of) a system is often just an intermediary step to other goals, such as fairness or appropriate reliance (Langer and Landers, 2021; Schoeffer et al., 2024). For example, by understanding how a particular system’s output came to be, prior work has argued that a person should be enabled to assess whether this output was based on valid criteria or not (Páez, 2019; Gilpin et al., 2019). If an unfavorable decision was based on (a proxy for) the skin color of a person, this person should supposedly be able to recognize that they were treated unfairly (Carvalho et al., 2019; Langer et al., 2021a).
@@
However, the goal of XAI is often narrowly construed as the development of methods to create interpretable surrogate models of black box models (Rudin, 2019; Dwivedi et al., 2023). Such a narrow view of XAI artificially restricts the space of options that can be chosen to facilitate stakeholders’ understanding of AI systems. Take, e.g., model cards for model reporting (Mitchell et al., 2019) or datasheets for data sets (Gebru et al., 2021). Although these approaches provide important information about certain aspects of an AI model (e.g., which training data was used) and thus improve stakeholders’ understanding of these aspects, they would rarely be counted as XAI. Further, XAI methods alone do not provide other kinds of highly relevant information, such as the normative motivations guiding the development of an AI model (Loi et al., 2021), the analysis of the social context in which it is deployed (Green, 2022), or the descriptive outcome statistics of the deployed model (Chouldechova, 2017).
@@
In a similar vein, the high hopes placed in XAI to mitigate issues of fairness often remain vague or unfulfilled (Deck et al., 2023). Prior works have criticized the explanatory value (Rudin, 2019; Herman, 2017), susceptibility to manipulations (Aïvodji et al., 2019; Lakkaraju and Bastani, 2020), and unsatisfactory interpretations (Balkir et al., 2022; Deck et al., 2023) of XAI methods. For example, popular feature-based explanations like LIME (Ribeiro et al., 2016) or SHAP (Lundberg and Su-In Lee, 2017) highlighting the use of sensitive features (e.g., gender and race) provides little information about fairness because these features are often correlated with proxy variables and embedded in use case-specific normative contexts (Dwork et al., 2012; Lipton et al., 2018; Nyarko et al., 2021). Instead, broadening the conception of XAI enables a more holistic and meaningful mapping of how information about an AI system can contribute to various fairness desiderata.
@@
Against this background, we propose to distinguish the narrow conception of XAI, solely pertaining to the inner workings of an AI system, from a broad conception of XAI that provides information beyond these inner workings and includes explanations of the broader socio-technical system (Dhanorkar et al., 2021). Specifically, we suggest that the broad conception of XAI includes all types of information that increases stakeholders’ understanding of (aspects of) an AI system. If not indicated otherwise, we will simply refer to XAI and mean the broad conception.
3. How XAI Can Be Leveraged for Fairness Along the AI Lifecycle

### v4

# EWAF’24: European Workshop on Algorithmic Fairness, July 01–03, 2024, Mainz, Germany

@@ Abstract
The widespread use of artificial intelligence (AI) systems across various domains is increasingly surfacing issues related to algorithmic fairness, especially in high-stakes scenarios. Thus, critical considerations of how fairness in AI systems might be improved—and what measures are available to aid this process—are overdue. Many researchers and policymakers see explainable AI (XAI) as a promising way to increase fairness in AI systems. However, there is a wide variety of XAI methods and fairness conceptions expressing different desiderata, and the precise connections between XAI and fairness remain largely nebulous. Besides, different measures to increase algorithmic fairness might be applicable at different points throughout an AI system’s lifecycle. Yet, there currently is no coherent mapping of fairness desiderata along the AI lifecycle. In this paper, we we distill eight fairness desiderata, map them along the AI lifecycle, and discuss how XAI could help address each of them. We hope to provide orientation for practical applications and to inspire XAI research specifically focused on these fairness desiderata.
keywords:
[…]
1 Introduction
@@
The emergence and widespread use of artificial intelligence (AI) systems across various sectors and domains is increasingly shifting attention from considerations of mere performance to considerations about algorithmic fairness. This is particularly relevant for systems employed in high-stakes scenarios and especially pressing in contexts prone to harmful societal biases [1, 2]. This has sparked a growing demand for approaches to scrutinize and improve fairness in AI systems. In the literature, there is a common recognition that fairness in AI systems demands various perspectives and measures (e.g., [3, 4, 5]). In this paper, we set out to integrate two strategies that have been suggested: First, a growing community of researchers proposes explainable AI (XAI) as a versatile and powerful tool to combat unfairness [6]. Second, others have focused on the AI lifecycle and tried to determine where fairness issues originate (e.g., [7, 8]). They hope that once identified, fairness issues can be mitigated by taking appropriate steps in the relevant phase within the lifecycle.
@@
While both these strategies seem intuitively promising, neither currently presents a satisfactory and comprehensive picture. A primary reason why neither approach currently fulfills their potential is, we take it, that there are various types and kinds of fairness discussed in the literature, yielding different fairness desiderata. Differentiating between these desiderata is crucial to gain an overall picture of which measures are most promising to address fairness in which contexts. Without such differentiation, the utility of XAI is not as straightforward as commonly claimed. Accordingly, there is a need for more clarity about how, exactly, XAI contributes to which fairness desideratum [6]. Existing discussions on this matter suffer, by and large, from both a too narrow conception of XAI and a mostly under-specified notion of fairness. Similarly, existing attempts to map measures for achieving fairness onto the AI lifecycle remain crucially incomplete, for they have limited their attention to only a subset of the relevant fairness desiderata. To overcome these limitations, we distill eight fairness desiderata, map them along the AI lifecycle, and discuss how XAI can help address each of them.
@@
For the purposes of this paper, we will focus specifically on how XAI can be utilized to improve AI systems’ fairness throughout their lifecycle. Concerning this question, we aim to develop a holistic account based on the fairness desiderata we distill. While our proposal is not intended as a guideline, we believe it will stimulate scientific discussions about and practical applications of fairness measures and, as such, will be a valuable starting point to explore fairness opportunities for researchers, developers, and regulators alike.
@@
We begin by introducing preliminaries and diagnosing the problem in Section 2. Importantly, we do not limit our understanding of algorithmic fairness to computational perspectives. In Section 3 we propose eight fairness desiderata from interdisciplinary literature: fairness understanding, data fairness, formal fairness, perceived fairness, fairness with human oversight, empowering fairness, long-term fairness, and informational fairness (contribution #1). We map our fairness desiderata along the different stages within the AI lifecycle and suggest that each fairness desideratum affords a different entry point for taking measures to improve fairness. This closes the first gap in the current literature as it highlights where in the lifecycle different fairness desiderata become especially relevant (contribution #2). Utilizing this mapping, we discuss how XAI can be leveraged to address the different fairness desiderata at the respective points throughout the AI lifecycle. This closes the second gap in the current literature as it allows us to systematically examine the potential of XAI to address algorithmic fairness in different circumstances (contribution #3). To illustrate the utility of our approach, we discuss its application to the COMPAS case (see [9]) in Section 4. Before closing, we point to some avenues for future research in Section 5.
2 Background
@@
In this article, we contribute to closing two research gaps: 1) considerations of fairness along the AI lifecycle are incomplete, and 2) it is unclear how exactly XAI can help foster fairness. Before closing these gaps, we retrace contributions and debates of prior works.
2.1 Algorithmic Fairness
@@
The debate on algorithmic fairness draws inspiration from various disciplines. Several scandals surrounding AI systems that disadvantage marginalized groups [9, 10, 11] have shattered early hopes put into the “neutrality” of AI [1]. In reaction, scholars from computer science, philosophy, social science, law, and psychology have been engaging in debates on what it means for algorithmic decision-making to be fair (see [12] for interdisciplinary perspectives).
@@
The technical debate on algorithmic fairness primarily focuses on formal measures of fairness [13, 14, 15]. After it has proven unhelpful to remove information about membership in marginalized groups from training data (“fairness through unawareness” [16]), most approaches from the field of computer science rely on comparing the outcomes for people from different groups. An important finding is that the different formal measures for fairness can, in most cases, not be fulfilled simultaneously [17, 18]. Thus, (formally) fair AI systems require crucial design choices about which formal measures to apply in which context.
@@
To aid such decisions, philosophers have connected formal measures to different philosophical theories [19, 20, 21]. In general, “fairness” is a normative concept—and debates about fairness can be understood as a way to discuss what is morally right or wrong in a given case (see [12]). More specifically, fairness can be understood as an issue of justice, non-discrimination, or equality [19], which connects it to numerous strands of philosophical discussion.
@@
Beyond that, social scientists have pointed out shortcomings of existing formal methods, e.g., problems in detecting issues of intersectionality [22] and their reliance on constructed categories like race and gender [23]. Legal scholars are primarily interested in how legislation on discrimination can be applied to algorithmic decision-making and whether additional regulation is required [24, 1]. Psychological research is generally interested in whether algorithmic decision-making is perceived as fair [4, 25]. For example, Colquitt and Rodell [26] distinguish four different psychological dimensions of fairness: Whether the outcome is fair (viz., distributive justice), whether the process that leads to the outcome is fair (viz., procedural justice), whether the information about the decision is communicated truthfully and thoroughly (viz., informational justice), and whether individuals are treated respectfully (viz., interpersonal fairness).
@@
Overall, algorithmic fairness can be addressed from several different angles—some of which are complementary, some of which are contradictory [27, 18]. This stresses the need to be clear about the specific desideratum behind any attempt to improve fairness.
2.2 Considerations of Fairness Along the AI Lifecycle Remain Incomplete
@@
A common strategy to determine potential entry points to address fairness issues involves examining each step of the AI lifecycle. There have been several proposals of prototypical AI lifecycles (e.g., [28, 29, 30, 31, 32]), each with slightly different stages or adapted to different application areas. The AI lifecycle used in this article is a combination of the proposals by Quemy [31] and Wang et al. [30] involving the following stages: 1) problem formulation, 2) data collection, 3) data analysis, 4) feature selection, 5) model construction, 6) model evaluation, 7) deployment, and 8) inference and usage (see Figure 1).
Figure 1: The AI Lifecycle we use, combined from Quemy [31] and Wang et al. [30].
@@
First, problem formulation involves abstraction and formulation of the problem, which is to be solved using AI. Afterward, data collection aims to establish a representative data set suitable to develop an AI system able to solve the defined problem. Data analysis includes descriptive statistics to understand the characteristics of the data at hand and data pre-processing to prepare the data for further operations. Subsequently, in the feature selection stage, features are excluded, transformed, or aggregated for effective and efficient handling in the training process. Model construction, then, includes the selection of the training algorithm and the training itself. Iteratively, performance objectives are tested during model evaluation and optimized through algorithm adjustments, e.g., parameter tuning. Deployment refers to the integration of the AI system into a productive environment once the model achieves sufficient performance. Finally, inference & usage describes concrete output generated by the AI system and its impact on the surroundings, such as the business context, society, and environment. Note that these stages are often not entered sequentially, leaving room for iterations and loops between stages [33].
@@
Much research has focused on what types of bias can emerge in or affect different steps of the AI lifecycle (e.g., [3, 8, 34, 35]). For example, Singh et al. [35] take the CRISP-DM model [28] and review the types of bias that can occur at different stages of this model. However, these considerations remain incomplete, as most of these works have a strong technical focus and deal only with formal notions of fairness, e.g., by providing guidance for choosing fairness metrics or debating the role of potential fairness-utility tradeoffs (e.g., Castelnovo et al. [36], De-Arteaga et al. [3]). Against this background, it is important to develop a view that also maps other conceptions of fairness into the lifecycle.
2.3 Considerations of XAI for Fairness Remain Incomplete
@@
To amend the lack of understanding of AI-based systems, their reasoning processes, and their outputs, the research field of XAI emerged in recent years [37, 38, 39, 40, 41]. Generally, XAI aims to provide a stakeholder with information about some aspect of an AI system to facilitate their understanding of this aspect [39]. Understanding (some aspect of) a system is often just an intermediary step to other goals, such as fairness or appropriate reliance [42, 43]. For example, by understanding how a particular system’s output came to be, prior work has argued that a person should be enabled to assess whether this output was based on valid criteria or not [38, 44]. If an unfavorable decision was based on (a proxy for) the skin color of a person, this person should supposedly be able to recognize that they were treated unfairly [45, 46].
@@
However, the goal of XAI is often narrowly construed as the development of methods to create interpretable surrogate models of black box models [47, 48]. Such a narrow view of XAI artificially restricts the space of options that can be chosen to facilitate stakeholders’ understanding of AI systems. Take, e.g., model cards for model reporting [49] or datasheets for data sets [50]. Although these approaches provide important information about certain aspects of an AI model [^1] Throughout this paper, we use the term “AI models” synonymously for “machine learning model” or “ML model”. While we acknowledge the technical distinction between AI and Machine Learning (ML) as discussed in [51], we adopt the use of the broader “AI” term as the prevalent terminology established in the scientific community. This choice reflects the contemporary linguistic trend rather than a lack of distinction between the two phenomena. (e.g., which training data was used) and thus improve stakeholders’ understanding of these aspects, they would rarely be counted as XAI. Further, XAI methods alone do not provide other kinds of highly relevant information, such as the normative motivations guiding the development of an AI model [52], the analysis of the social context in which it is deployed [53], or the descriptive outcome statistics of the deployed model [54].
@@
In a similar vein, the high hopes placed in XAI to mitigate issues of fairness often remain vague or unfulfilled [6]. Prior works have criticized the explanatory value [47, 55], susceptibility to manipulations [56, 57], and unsatisfactory interpretations [58, 6] of XAI methods. For example, popular feature-based explanations like LIME [59] or SHAP [60] highlighting the use of sensitive features (e.g., gender and race) provide little information about fairness. This is because these features are often correlated with proxy variables and embedded in use case-specific normative contexts [16, 61, 62].
@@
Instead, broadening the conception of XAI enables a more holistic and meaningful mapping of how information about an AI system can contribute to various fairness desiderata. Against this background, we propose to distinguish the narrow conception of XAI, solely pertaining to the inner workings of an AI system, from a broad conception of XAI that provides information beyond these inner workings and includes explanations of the broader socio-technical system [63]. Specifically, we suggest that the broad conception of XAI includes all types of information that increases stakeholders’ understanding of (aspects of) an AI system. For the remainder of this paper, if not indicated otherwise, we refer to the broad conception when we mention XAI.
3 How XAI Can Be Leveraged for Fairness Along the AI Lifecycle